5)Solve the integral equation f(t)=t+3 ∫0^t f(τ)dτ?

Explain the problem with step by step process?


1
Expert's answer
2021-10-03T20:32:02-0400

Considering the theorem: we have L[e 2t t 2 ] = F(s − 2) where L[t 2 ] = 2! s 3 = F(s), s > 0. Thus, L[e 2t t 2 ] = 2 (s−2)3 , s > 2. (b) As in part (a), we have L[e 3t cos 2t] = F(s−3) where L[cos 2t] = F(s−3). But L[cos 2t] = s s 2+4 , s > 0. Thus,



Since L[t] = 1 s 2 , by Theorem , we have



Therefore, our solution becomes: